A buffer solution contains 0.441 M ammonium bromide and 0.383 M ammonia. If 0.0318 moles of potassium hydroxide are added to 250 mL of this buffer, what is the pH of the resulting solution ? (Assume that the volume does not change upon adding potassium hydroxide) pH =
Answer #1

Given buffer contain two components. Acidic component is ammonium bromide and basic component is ammonia.

When base is added to buffer solution it reacts with acidic component and produces base. The reaction is

NH4Br + KOH KBr + NH3 + H2O

[KOH ] = NO of moles / Volume of solution in L

= 0.0318 mol / 0.25 L

= 0.1272 M

Let's use ICE table

Concentration NH4Br KOH NH3
Initial 0.441 0.1272 0.383
Change -0.1272 -0.1272 +0.1272
Equilibrium 0.3138 0.0000 0.5102

pH of buffer solution is calculated by using Henderson's equation.

pH = pKa + log [ NH3 ] / [NH4Br]

= 9.24 + log 0.5102 / 0.3138

= 9.24 + 0.211

= 9.45

ANSWER :pH = 9.45